RE: Dev Status - 20 August 2007

From: Daniel Lee <dlee_at_mediadefender.com>
Date: Tue, 21 Aug 2007 13:18:34 -0700

Nope, I didn't get this.

I talked to both Serge and Linus today. The QA team ran some tests for
Serge Friday and Monday. Serge discovered a problem with his ProxMaster
stuff--he says he has to fix it before we can do more testing. He's
working on things for MiiVi so he didn't give me an ETA on the fix. I
was going to ask him about it again tomorrow.

I spoke with Linus about Soulseek, and we had him change the usernames
for his protection to username+username length % 7 instead of mod 5
(Easier to identify his stuff vs. Dylan's.) We're going to run some
tests today. So far it looks to be hitting on some tracks and very
minimally on others.

-----Original Message-----
From: Ben Grodsky
Sent: Tuesday, August 21, 2007 1:12 PM
To: Daniel Lee
Cc: Jay Mairs
Subject: Fw: Dev Status - 20 August 2007

Did you get this yesterday?

Sergio says he's waiting to hear back from qa about prox master and
Linus should be asking you to start testing his soulseek protection
soon.

----- Original Message -----
From: Ben Grodsky
To: development
Sent: Mon Aug 20 12:34:20 2007
Subject: Dev Status - 20 August 2007

It's been a while since I've sent one of these. Jay's trying to
re-focus me on development, so I should be sending these out more
regularly after the weekly status meetings. These notes are meant to
help you remember what was discussed during the meeting: both your own
project(s); and your fellow team members' projects.

Sergio

* proxmaster - have something for QA to test

* java console

        * integrating http get for podcast downloading

        * web services already setup for using URLs as sources
(from Dylan)

        * atty general

                * needs to coordinate with Ivan mysql encryption
connection

                * will get passed parameters from Ivan.

                * Just give certificate path for connection

dylan

* miivi

        * podcast

                * filters out anything w/o thumbnail

                * if no channel thumb and no episode thumb, then
skipped

                * currently 1500 videos in system from Colin's
last list version (20-40 podcast feeds)

                * timer every ?? hr indexes

                * some feeds have url redirect and have to deal
with this condition (they create black thumbnails)

                * working on recategorize tool for Colin

        * soulseek

        * logging on, but didn't see own searches

        * so no update there

        * ed2k - no update

ebert

* data collection moving along well

* new IPs for BT (not being used on mininova/tpb b/c of active
moderating)

* meeting w/ Dr Hahn Thursday & Friday to discuss his findings
thus far

* postmaster banned from ticketing system (may resolve mailbombing
issue)

gerald

* added another site to searcher

* fixed problem for .torrent downloading

* migrating db file

simran

* shareaza

        * problem passing msgs between hubs and leaves

        * changed code to do buffer checks ???

        * socket times out from leaf

                * regular shareaza works - so IP doesn't sound
banned

                * possibly mistake in handshake - checking in
wireshark

                * no current stability problems with jeff's legacy
system

danny

* UMG data

        * storing filenames and track number and dumped a couple
weeks of data

        * still processing data automatically for Ares (9-10 hrs
to process 1 day's data); some issue with Kad has it behind; gnutella is
almost caught up

                * let Jay know if it doesn't seem the data
processing will catch up to that day's data

                * may want to split up data processing to 2+
computers

                * set up another computer now (get from Ivan) that
can start processing data from yesterday. We should have 2 systems
working on ares data processing: one catching up and one working on
current data.

        * downloading

        * getting 10% of new hashes

        * similar to discussions with Ivan, Sujay and Abishek,
download by hashes with most sources first, but keep track of failed
download attempts so you can re-try those later.

        * for now, follow Ivan's format of directory structure how
to upload to UMG ftp (filename = hash.extension; real filename from p2p
will be in database)

jed

* inflation load balanced now for multiple computers

* removed redundancies from Nutter's BT scraped torrent files db
(57mm --> 7mm rows)

* redid AP (removing Nutter code and replacing with new Jed code)

Andrew

* mi.mediadefender.com

        * added logout button to site

        * adding projects to system

        * adding functionality to pass-off project input system to
PM team

        * working on errors in geo IP data

        * added Universal tracks (waiting to activate from Sujay's
ok - he's adding more machines)

sheetal

* finished reactiontorrents tracker

linus

* moving to QA

* worked with Sujay to optimize

abishek

* UMG downloader

        * try to redownload hashes that may have accumulated more
sources over time

        * track how many successful downloads out of hashes tried,
so that we can figure out whether more system need to be added

        * working on optimizations with Sujay

        * successfully downloaded 500+ hashes (sorted by sources)
out of data set of ~3000 hashes (with >1 source)

        * for now, follow Ivan's format of directory structure how
to upload to UMG ftp (filename = hash.extension; real filename from p2p
will be in database)

sujay

* some servers on DCa aren't working right, fixing those

* made sponsor link change

ivan

* atty general

        * has to coordinate with Sergio mysql encryption for atty
general investigation

        * will test downloader on local computer and then let Jay
know to move-along w/ atty general

        * umg ares downloader

        * trying to download everything newly found that day

        * currently downloading a few 100 out of the several
thousand found

        * just have to change u/l searcher

        * piolet

        * looking into mysql encryption with Sergio
Received on Fri Sep 14 2007 - 10:55:51 BST

This archive was generated by hypermail 2.2.0 : Sun Sep 16 2007 - 22:19:45 BST